The position (x)of a particle on a straight line motion is given by x=2=10t-5tsquare(m).is velocity(v)is best represented by

Grade:12th pass

2 Answers

Khimraj
6 years ago
The position (x)of a particle on a straight line motion is given by x=2-10t-5tsquare(m).
the velocity of a particle is given by v = dx/dt.
SO v = 10 – 10t.
Hope it clears.
Khimraj
6 years ago
The position (x)of a particle on a straight line motion is given by x=2-10t-5tsquare(m).
the velocity of a particle is given by v = dx/dt.
SO v = -10 – 10t.
Hope it clears.